One day Daniel Bernoulli was sitting in a Starbucks sipping on a Venti Mocha Frappuccino. He pondered some of the many applications of his newly developed principle. Grab a Starbucks of your own, and help him to answer the following questions. Assume air wehs 0.081 lb/ft3 NOTES IMAGES DISCUSS UNITS STATS HELP Part Description Answer Chk History In a hurricane, roofs of structures are really lifted off, and not blown off, due to the decreased pressure as the wind blows over the roof. Determine the required wind speed to lift a roof that welghs 17 psf Assume the roof is not tied down (include units with answer) 18.33 pts. 110% 2%try puraty Hints: 1.0 # tries: 0 Show Details Fomat Cheok Determine the minimum required take-orr speed of a 29000 pound UScalrways jet with a wing area of 1420 ft2, Use a lift coefficient of K-0.12 (include units with answer) 18.33 pts, 1 10% 2 try peraty Hints: 1.0 B. # tnes : O Snow Details Format Ghook Daniels good friend Giovanni Venturi joined him for another round of fraps. They were discussing ways to determine the velocity of water in a pipe. If the pressure drops 390 kPa as water flowing in a 50 cm diameter pipe is forced to flow through a 20 cm diameter portion, what is the velocity of the water in the 50 cm section of pipe? (include units with answer) 18.33 pts,110% C. # tries: ( Snow Details 2% try par.tv Hints: 10
Answer #1

A) Here, we are using the following formula for v to calculate speed v

ひ 2P

P is the pressure, rho is the air density and putting all the parameters in SI units

2 x 813.9835.43m/s V) 1.297

B) Here, two forces are acting one is the lifting force F which can be given as

F=rac{1}{2} ho v^2K A......(1)

rho is the air density, v is the speed that needs to be determined and K, A are lifting coefficient and area, respectively. The second force f2 is given as

2mg.. 2)

Lifting will take place only if two forces balancing each other

rac{1}{2} ho v^2K A=mg

Solving for v, we get

mg 1 V)

Here, we are putting all the values in SI system

2 × 1315.4 × 9.8 1.297× 0.12 × 131.9 17.35m/s V)

(c) Here, we need to use the Bernoulli;s principle and equations which can be given as follows

A_1 v_1=A_2 v_2.....(1c)

P_1+rac{1}{2} ho v_1^2+ ho gh=P_2+rac{1}{2} ho v_2^2+ ho gh.....(2c)

We have used the subscripts 1 for 50 cm diameter pipe and subscript 2 for 20 cm diameter pipe. A, v, P, g, h represent area, speed, Pressure, acceleration due to gravity and height, respectively... We have assumed that the height h remains the same. The cross sectional area is circular and is equal to pi r2 with r is the radius. Firstly, we solve the first equation to find v2 in terms of v1 as we are looking to find v1/

T2 1

(25)21,1-(10)21,2

v_2=6.52 v_1......(3c)

Pressure drop P2-P1=390KPa so using Eq. (2c) along with Eq. (3c) to find v1

P_2-P_1=rac{1}{2} ho (v_1^2-v_2^2)

2 390 × 10°- 1.29ให้ (1-6.529

-390 x 103-26.92v

1,1-120.36m/s